Can someone explain how I am supposed to solve by substitution with the problems x-y=5 x+2y=7 and 3b+2a=2 -2b+2a=8
Answer by jai_kos(139)

pROBLEM 1

x-y=5

x+2y=7
sOLUTION:
x-y = 5
==> x = 5 + y

Now substitute the value x = 5 + y in the second equation x + 2y =7

==> 5 +y + 2y = 7

==> 5 + 3 y = 7

==> 3y = 7 -5
==> y = 2/3

Now again we have x =5 + y

==> x = 5 + 2/3

==> x = 17 /3

so the solution is x = 17/3 and y = 2/3

----------------------------------------------------------------------------
3b+2a=2

-2b+2a=8

==> 2a = 8 + 2b

==> a = 8/2+ 2b/2

==> a = 4 + b

now substitute the value a = 4+b in the first equation, 3b + 2a = 8

==> 3b+ 2 ( 4 + b ) = 8

==> 3b+ 8 + 2b = 8

==> 5b = 8-8

==> 5b = 0

==> b = 0/5

==> b = 0

Now we have, a = 4 + b

==> a = 4 + 0

==> a = 4.

therefore solution is a = 4 and b = 0